OpenStudy (anonymous):

what is the derivative of the square root of 6x+10

OpenStudy (anonymous):

\[\huge (6x+10)^{1/2} = \color{green}{ \frac{1}{2}(6x+10)*6}\]

OpenStudy (anonymous):

\[\huge \color{green}{ dx = 3*(6x+10)^{-1/2}}\]

OpenStudy (anonymous):

chain rule

OpenStudy (anonymous):

thanks do you happen to know what the derivative of e^ln(x) is as well?

OpenStudy (anonymous):

chain rule again

OpenStudy (anonymous):

so e^x is one functions and ln(x) is another?

OpenStudy (anonymous):

not quite you can make it easy by eliminating e^ln since e raise to natural log = 1 so e^ln(x) would equal to x now find the derivative of x

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem